OpenStudy (rainbow_rocks03):

PLZ help me ASAP i will medal someone for best response.QUESTION:A piece of tile is 4 1/2 inches long and 3 1/4 inches wide. What is the area of the tile? A.7 3/4 B.12 1/8 C.12 D.14 5/8

OpenStudy (anonymous):

what do you think it is

OpenStudy (rainbow_rocks03):

thx @Juando

OpenStudy (anonymous):

if you have a math question just go to mathway

OpenStudy (anonymous):

No if you have a math question you should ask for help... not the answer.

OpenStudy (anonymous):

\[4\frac{ 1 }{ 2 }=4+\frac{ 1 }{ 2 }=\frac{ 8+1 }{ 2 }=\frac{ 9 }{ 2 }\] \[3\frac{ 1 }{ 4 }=3+\frac{ 1 }{ 4 }=\frac{ 12+1 }{ 4 }=\frac{ 13 }{ 4 }\] area of tile\[=\frac{ 9 }{ 2 }\times \frac{ 13 }{ 4 }=\frac{ 117 }{ 8 }=14\frac{ 5 }{ 8 }\]